OpenStudy (anonymous):

Hello, I need help with factoring quadratics: 64a2 - 48a + 9 I only know how to factor ones with no coefficient, please help!

OpenStudy (kainui):

You could look at it like: \[64(a^2-\frac{48}{64}a+\frac{9}{64})\] and solve that one in parenthesis then multiply it by 64 afterwards.

hero (hero):

Hint 24 + 24 = 48 24*24 = 576

hero (hero):

64a^2 - (24 + 24)a + 9 64a^2 -24a - 24a + 9

hero (hero):

8a(8a - 3) -3(8a - 3) @OceanBEAUTI16 are you following?

OpenStudy (anonymous):

So how are you getting the eight?

hero (hero):

64a^2 - 24a = 8a(8a - 3) -24a + 9 = -3(8a - 3) By factoring out the GCF of each set of binomials.

OpenStudy (anonymous):

Thanks! How about this one, since it doesn't have a middle term. 16n^2 - 49

hero (hero):

For that use the difference of squares rule: \[a^2 - b^2 = (a + b)(a - b)\]

OpenStudy (anonymous):

So that's what it means, thank you :)

ganeshie8 (ganeshie8):

these all are special quadratics, u should memorize below identities : \(\large (a+b)^2 = a^2+2ab + b^2\) \(\large (a-b)^2 = a^2-2ab + b^2\) \(\large a^2-b^2 = a^2-b^2\)

hero (hero):

In this case a = 4n b = 7 And a^2 = (4n)^2 = 16n^2 b^2 = (7)^2 = 49

OpenStudy (anonymous):

I love squares, they're so much easier.

hero (hero):

64a^2 - 48a + 9 = (8a - 3)^2 Which is also a square

OpenStudy (anonymous):

Right! Sometimes things just go over my head *ahem*. Thanks!